A capacitor consists of two closely spaced metal conductors of large area, separated by a thin insulating foil. It has an electrical capacity of 3200.0 μF and is charged to a potential difference of 84.0 V. Calculate the amount of energy stored in the capacitor. Tries 0/20 Calculate the charge on this capacitor when the electrical energy stored in the capacitor is 9.73 J. Tries 0/20 If the two plates of the capacitor have their separation increased by a factor of 2 while the charge on the plates remains constant, by what factor is the energy stored in the capacitor increased?
Answer #1

Solution Given : C = 3200.00μ F = 3200.00 × 10-6F V = 84.0V

Now Energy stored on the capacitor by following formula Plugging in values in the above equation, we get E = ã × (3200.00 × 1